OpenStudy (anonymous):

use Gauss's approach to find the follwoing sums (do not use formulas). a... 1+2+3+4+...+999 b... 1+3+5+7+...+997 the sum of sequance a is... the sum of sequance b is...

OpenStudy (amistre64):

is there a reason they want you to add up the same number 1000 times?

OpenStudy (anonymous):

no. that's just what is it asking and I have no clue how to do it!

OpenStudy (amistre64):

without a formula, then the only alternative would be by hand ....

OpenStudy (anonymous):

how would you do it??

OpenStudy (amistre64):

by using the formula

OpenStudy (dumbcow):

haha ..i thought Gauss's approach was the formula ? he came up with the formula..allegedly

OpenStudy (amistre64):

spose we do this on a smaller set: {1,2,3,4,5} add the set to itself, then we have twice as much in value. {1,2,3,4,5} + {1,2,3,4,5} = 2{1,2,3,4,5}

OpenStudy (amistre64):

if we add the set to itself in a convienent way we get: {1,2,3,4,5} {5,4,3,2,1} ---------- {6,6,6,6,6} ; or simply 5(6) but this amount is twice the one we want, so divide it in half

OpenStudy (amistre64):

why you would want to do that for 1000 terms is beyond me

OpenStudy (anonymous):

I did that for my problem, and it was wrong.

OpenStudy (amistre64):

its just that simple for part a, for part b you have to determine how many terms there are ... there aint 997 of them

OpenStudy (dumbcow):

\[\text \sum = \frac{n(a_1 +a_n)}{2}\]

OpenStudy (amistre64):

in part b, if we include the evens to fill out the set; we get from 1 to 998; we only want half of those terms; so n = 998/2 in that case

OpenStudy (dumbcow):

Gauss noticed that if you add the first and last number, and then the 2nd and 2nd to last number, you get same number 1+999 = 1000 2+998 = 1000 ... the number of pairs is 999/2 --> (999/2) *1000

OpenStudy (anonymous):

nope. just tried that and it's wrong.

OpenStudy (dumbcow):

im pretty sure its right sum = 499,500

OpenStudy (anonymous):

The answer was 250,000. How.. I have no idea. I am so lost!!

OpenStudy (anonymous):

Try this one... 47+48+49+50...+134

OpenStudy (amistre64):

\[\frac{(134-46)(47+134)}{2}\]

OpenStudy (anonymous):

Thank you.. that's right. but so I understand, why use 46 and not 48, 49, or 50?

OpenStudy (amistre64):

it starts at 47, its simpler to count the number of terms when they start at 1; 47-46 = 1 subtract the same amount from both ends

OpenStudy (amistre64):

47 to 134 -46 -46 --- ---- 1 to 88 ; therefore there are 88 terms in the given sequence

OpenStudy (amistre64):

of course im taking notice that each term is "1" from the next.

OpenStudy (anonymous):

ok.. i think i understand.. sort of